OpenStudy (anonymous):

help please! The table below shows the square roots of different numbers: Number (x) 4 4 9 9 Square root (y) 2 -2 3 -3 Part A: Does the table represent y as a function of x? Justify your answer. (5 points) Part B: The total cost f(x), in dollars, for renting a car for a week and driving it x miles is shown below: f(x) = 90 + 0.11x What is the value of f(300), and what does f(300) represent? (5 points)

OpenStudy (michele_laino):

hint: the point x=4 has two images, namely +2 and -2, similarly for point x=9 what can you conclude?

OpenStudy (anonymous):

it is not a function?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

oh so that would be the answer to part a!? thats so much simpler than i thought

OpenStudy (anonymous):

now i don't understand where f(300) comes into play?

OpenStudy (michele_laino):

yes! since your function, doesn't check the general definition of a function

OpenStudy (michele_laino):

it is simple, we have to replace x with 300, so we have to compute this: \[\Large f\left( {300} \right) = 90 + 0.11 \cdot 300 = ...?\]

OpenStudy (anonymous):

oh so that would equal 123?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

but what does it represent?

OpenStudy (michele_laino):

it is represent the image of the point x=300, by means of your function: 

OpenStudy (michele_laino):

oops..it represents...

OpenStudy (anonymous):

ooh i think i see, alright thank you so much for your help! @Michele_Laino

OpenStudy (michele_laino):

:)